OpenStudy (anonymous):

The height of a triangle is 4 more than the base, and the area of the triangle is 6 square units. Find the length of the base. Let x = the length of the base.

OpenStudy (anonymous):

x........length of the base 4x........height area=1/2*base*height 1/2*4x^2=6 2x^2=6 x^2=3 x=square root of 3...